OpenStudy (anonymous):

sinxsecxtanxcscx = sinx/1-sin^2x

OpenStudy (kainui):

When you see nonsense like this, always, always, always turn everything into sines and cosines. So for instance, if you see tanx write sinx/cosx instead.

OpenStudy (anonymous):

yup\[sinx \times 1/cosx \times sinx/cosx \times 1/sinx\]

OpenStudy (anonymous):

cancelling those out \[sinx/\cos^2x\]

OpenStudy (kainui):

Let him do it!

OpenStudy (anonymous):

cos^2x + sin^2x = 1 then cos^2x = 1-sin^2x

OpenStudy (anonymous):

then you get sinx / (1-sin^2x) tada!

OpenStudy (bibby):

tada! it's magic! get someone to do your homework for you! INTERNET

OpenStudy (anonymous):

am I make myself clear

OpenStudy (kainui):

"How to get through school without doing anything" written and illustrated by: the internet.

OpenStudy (anonymous):

@Kainui yes, will be the bestseller book

OpenStudy (kainui):

Sadly.

OpenStudy (anonymous):

tanx can be return as sinx/cosx

OpenStudy (anonymous):

secx can be written i/cosx,then we get sinx/cos^2x and cos^2x can be written as 1-sin^2xthen we sinx/1-cos^2 x